# KB-DMGen: Knowledge-Based Global Guidance and Dynamic Pose Masking for Human Image Generation

### v1

@@ Abstract
Recent methods using diffusion models have made significant progress in human image generation with various control signals such as pose priors. In portrait generation, both the accuracy of human pose and the overall visual quality are crucial for realistic synthesis. Most existing methods focus on controlling the accuracy of generated poses, but ignore the quality assurance of the entire image. In order to ensure the global image quality and pose accuracy, we propose Knowledge-Based Global Guidance and Dynamic pose Masking for human image Generation (KB-DMGen). The Knowledge Base (KB) is designed not only to enhance pose accuracy but also to leverage image feature information to maintain overall image quality. Dynamic Masking (DM) dynamically adjusts the importance of pose-related regions. Experiments demonstrate the effectiveness of our model, achieving new state-of-the-art results in terms of AP and CAP on the HumanArt dataset. The code will be made publicly available.
@@
Figure 1: Overview of our method. The visual Knowledge Base (KB) provides global guidance, while the Dynamic pose Mask (DM) enables fine-grained local control. Together, they achieve unified control of both global semantics and local details.
@@ I Introduction
Previous methods [7, 8, 9] require a source image during training using variational autoencoders (VAEs) [10] or Generative Adversarial Networks (GANs) [11] for dictating the style of the generated images. These methods synthesize target images with specific human features by adjusting the source images, but the training process of these methods is unstable and highly dependent on the distribution of the source images. In contrast, recent advances in controllable text-to-image (T2I) Stable Diffusion (SD) [12] show potential to eliminate the need for source images, enabling greater creative freedom through reliance on text prompts and external conditions [3, 13, 4, 14].These methods often face challenges in accurately matching conditional images with sparse representations such as skeleton pose data [2]. To achieve accurate pose control, various pose guided T2I methods are proposed, such as introducing pose heatmap supervision loss [1], establishing a graph topological structure between the pose priors and latent representation of diffusion models [15], and applying pose masks [2]. These strategies effectively guide the network to focus on pose regions, thereby improving pose fidelity.
While precise pose alignment is essential, high-quality human image generation also demands coherent global structure and semantics. However, these methods [15, 1, 2] emphasize the modeling of pose details while neglecting global structure and semantics, which can easily lead to a decline in the overall quality of the images. To address this issue, we propose Knowledge-Based Global Guidance and Dynamic Masking for Human Image Generation (KB-DMGen). It introduces a visual knowledge base to provide global semantic guidance during image generation, maintaining image quality while improving pose accuracy. Meanwhile, the dynamic masking mechanism adaptively adjusts the weights of pose-related regions, enabling the model to better balance local pose precision. In summary, the contributions of this paper are as follows:
  • Designing a visual knowledge base to improve pose generation accuracy while preserving image quality.
  • Designing dynamic masks to control the precise generation of poses.
@@
Figure 3: (a) The codebook training. (b) The classifier training .
@@ II Related work
Pose-Guided Human Image Generation. Previous pose-guided approaches take a source image and pose as input to generate photo-realistic human images with preserved appearance, relying on GANs/VAEs for conditional generation [7, 9, 8, 16, 17]. Recently, Bhunia et al. [18] introduce texture diffusion modules and disentangled classifier-free guidance to accurately model appearance-pose relationships and ensure input-output consistency. Zhang et al. [19] propose a novel Dual-task Pose Transformer Network (DPTN), which introduces an auxiliary task (i.e., source-tosource task) and exploits the dual-task correlation to promote the performance of PGPIG. Shen et al. [20] presented a Progressive Conditional Diffusion Model (PCDM), which narrows the gap between character images under target poses and source poses step by step via a three-stage procedure. The above methods [18, 19, 20] remain highly dependent on the distribution of source images, operating under the paradigm of pose transfer that necessitates original image inputs. In contrast, recent work based on pose guided T2I dispenses with the need for source images entirely: HumanSD [1] enhances pose accuracy via heatmap-guided losses and StablePose [2] employs coarse-to-fine masking for precise alignment. GRPose [15] establish a graph topological structure between the pose priors and latent representation of diffusion models to capture the intrinsic associations between different pose parts. However, these methods [1, 2, 15] prioritize pose fidelity and lack global image quality assurance, prompting us to introduce a visual knowledge base for holistic realism and a dynamic mask for pose precision.
@@
Relation to VQVAE [28]. VQ-VAE [28, 29, 30] introduced the concept of learning discrete latent representations by mapping continuous features into a codebook of quantized vectors. This foundational framework paved the way for subsequent advancements in generating structured, discrete token spaces for various tasks. VQVAE2 [31] uses Vector Quantized Variational AutoEncoder (VQVAE) for large scale image generation. However, the pixel-wise sampling by auto-regressive models is time-consuming. Building on this, Zero-shot T2I [21] explores using a VQVAE-based framework to generate images from text descriptions. Text2Human [32] also uses a VQVAE-style codebook for modeling local clothing attributes to in human image generation. In contrast, we focus on aligning language and vision through a shared discrete token space. Specifically, image features are discretized to form a visual semantic codebook, and a classifier is trained to retrieve relevant tokens based on textual prompts, providing more efficient and interpretable semantic guidance for image generation.
@@ III Method
This section presents the overall framework, followed by detailed training of the codebook and classifier in the knowledge base, and concludes with the dynamic pose masking strategy.
@@ III-A Overview of KB-DMGen
Our goal is to generate high-quality human images conditioned on pose priors. As shown in Fig. 2, our proposed KB-DMGen framework integrates a Diffusion Pipeline, a Knowledge-Based Adapter, and a Dynamic pose Mask Adapter. The language-driven KB Adapter improves semantic alignment and image quality, while the DM Adapter enhances pose accuracy through refined structural guidance.
We adopt Stable Diffusion [12] as our base model, which includes a Variational Autoencoder (VAE) with encoder $E$ and decoder $D$, and a U-Net for denoising and image refinement in the latent space. Given a pose condition $c_{p}\in\mathbb{R}^{H\times W\times C}$ and a text prompt, we use a CLIP text encoder [33] to extract the text embedding $c_{t}$.
The input image is first mapped into a latent representation $z_{0}$ via the VAE encoder. The diffusion process is then guided by the text and pose conditions to iteratively denoise $z_{t}$ using the U-Net, aiming to approximate the data distribution $p(x)$ through maximizing the log-likelihood:
$\mathcal{L}_{d}=\mathbb{E}_{\mathbf{z}_{t},\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\left\|\epsilon_{t}-\epsilon_{\theta}(\mathbf{z}_{t},\mathbf{c}_{t},\mathbf{c}_{p},t)\right\|^{2}\right]$ (1)
where $z_{t}$ denotes the latent representation at time step $t$, $\mathcal{N}$ denotes the standard normal distribution, $\epsilon_{\theta}$ denotes the noise prediction network and $\epsilon_{t}$ denotes real noise.
The KB-DMGen adapter is a trainable module (KB Frozen) consisting of two components. The Text-query KB extracts visual codebook features from the KB that align with the text prompt, guiding image generation. The DM transforms pose into an attention mask that dynamically adjusts with timesteps. Together, these components guide the diffusion model to generate high-quality images $\hat{x}$.
@@ III-B Knowledge Base
Codebook training The goal of Codebook Training is to learn a discrete representation of image features that can be queried by text. The codebook consists of trainable entries $e={e_{1},e_{2},\dots,e_{K}}$, where $e\in\mathbb{R}^{K\times C}$, with $K$ denoting the number of codebook entries and $C$ the feature dimension.
@@
Given an image, we encode it using CLIP to obtain dense image features $z_{e}\in\mathbb{R}^{N\times C}$, where $N=H\times W$ is the number of spatial tokens and $C$ is the feature dimension. We then compute the squared Euclidean distance between each token in $z_{e}$ and the codebook entries:
$\mathbf{D_{\{i,j\}}}=\|z^{(i)}_{e}\|^{2}+\|e_{j}\|^{2}-2z^{(i)}_{e}e_{j}^{T},$ (2)
where $z_{e}^{(i)}\in\mathbb{R}^{C}$ is the $i$-th image token feature, $e_{j}\in\mathbb{R}^{C}$ is the $j$-th codebook entry, and $\text{sim}_{i,j}$ denotes the squared distance between token $i$ and codebook entry $j$.
This computation results in a similarity matrix $D_{\{i,j\}}\in\mathbb{R}^{N\times K}$, where each row corresponds to an image token and each column to a codebook entry. We then find the nearest codebook entry for each token via:
$a_{i}=\arg\min_{j}\mathbf{D}_{i,j},\quad i=1,\ldots,N,$ (3)
resulting in an index vector $a\in\mathbb{R}^{N}$, where $j\in\{1,2,\cdots,K\}$ and $a_{i}$ represents the codebook index that best matches the $i$th image position element indicates the nearest codebook vector. We then convert the indices $a$ into a one-hot encoding matrix $O\in\mathbb{R}^{N\times K}$, where:
$O_{ij}=\begin{cases}1&\text{if }j=a_{i},\\ 0&\text{otherwise}.\end{cases}$ (4)
Finally, the quantized feature representation $\mathbf{z}_{q}\in\mathbb{R}^{N\times C}$ is obtained by a matrix multiplication between the one-hot matrix and the codebook:
$z_{q}=O\cdot e,$ (5)
where $z_{q}\in\mathbb{R}^{N\times C}$ and $\cdot$ is matrix multiplication. To train the codebook, we use the reconstruction loss:
$\mathcal{L}_{\text{mse}}=\text{MSE}(z_{q},z_{e}).$ (6)
@@
Classifier training The second stage of knowledge base construction trains a classifier to align text features with the learned discrete visual tokens (codebook entries). Given an input text, we extract a global embedding $z_{t}\in\mathbb{R}^{C}$ using a pretrained text encoder (e.g., CLIP). The text embedding $t\in\mathbb{R}^{C}$ is mapped to token logits via a projection:
$y=\text{MLP}(z_{t}),$ (7)
where $y\in\mathbb{R}^{N\times K}$. We supervise the prediction using the discrete indices $a$ obtained from codebook training, optimizing the cross-entropy loss:
$\mathcal{L}_{\text{cls}}=\text{CE}(y,a).$ (8)
This learning process encourages the alignment of textual features with the corresponding visual codebook entries, thus enabling text-conditioned visual token retrieval.
III-C Dynamic Masking
To guide the attention mechanism towards human body regions while suppressing background interference, we introduce a dynamic masking strategy. A binary mask $m\in\{0,1\}^{B\times H\times W}$ is derived from the input pose image, where 1 denotes foreground (human) and 0 indicates background.
Given a diffusion timestep embedding $e\in\mathbb{R}^{B\times C}$, we apply a lightweight MLP to generate a soft gating value:
$g=\text{MLP}(e)\in(0,1).$ (9)
We construct a soft mask $\tilde{m}$ by combining the binary mask with the gating value:
$\tilde{m}=(1+g)\cdot m,$ (10)
@@ III-C Dynamic Masking
This soft mask is directly applied to the attention logits via element-wise multiplication (Hadamard product, denoted by $\odot$) before the softmax:
$\text{Attn}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\odot\tilde{m}\right)V,$ (11)
where $\odot$ indicates element-wise multiplication between the attention map and the soft mask $\tilde{m}$. It enables dynamic modulation of pose attention weights by adjusting the mask weights over diffusion timesteps.
@@
KB-DMGen 51.71 71.40 1.53 10.29 2.45 32.45
Figure 4: The visualization results of KB-DMGen.

### v2

# KB-DMGen: Knowlege-Based Global Guidance and Dynamic Pose Masking for Human Image Generation

@@ Abstract
Recent methods using diffusion models have made significant progress in Human Image Generation (HIG) with various control signals such as pose priors. In HIG, both accurate human poses and coherent visual quality are crucial for image generation. However, most existing methods mainly focus on pose accuracy while neglecting overall image quality, often improving pose alignment at the cost of image quality. To address this, we propose Knowledge-Based Global Guidance and Dynamic pose Masking for human image Generation (KB-DMGen). The Knowledge Base (KB), implemented as a visual codebook, provides coarse, global guidance based on input text-related visual features, improving pose accuracy while maintaining image quality, while the Dynamic pose Mask (DM) offers fine-grained local control to enhance precise pose accuracy. By injecting KB and DM at different stages of the diffusion process, our framework enhances pose accuracy through both global and local control without compromising image quality. Experiments demonstrate the effectiveness of KB-DMGen, achieving new state-of-the-art results in terms of AP and CAP on the HumanArt dataset. The project page and code are available at https://lushbng.github.io/KBDMGen.
@@
Figure 1: Overview of our method. The visual Knowledge Base (KB) provides global guidance by encoding visual features related to the text description, while the Dynamic pose Mask (DM) enables fine-grained local control. Together, they achieve unified control of both global semantics and local details.
@@ I Introduction
Previous methods [7, 8, 9] require a source image during training using variational autoencoders (VAEs) [10] or Generative Adversarial Networks (GANs) [11] for dictating the style of the generated images. These methods synthesize target images with specific human features by adjusting the source images, but the training process of these methods is unstable and highly dependent on the distribution of the source images. In contrast, recent advances in controllable text-to-image (T2I) Stable Diffusion (SD) [12] show potential to eliminate the need for source images, enabling greater creative freedom through reliance on text prompts and external conditions [3, 13, 4, 14]. These methods often face challenges in accurately matching conditional images with sparse representations such as skeleton pose data [1]. To achieve accurate pose control, various pose guided T2I methods are proposed, such as introducing pose heatmap supervision loss [1], establishing a graph topological structure between the pose priors and latent representation of diffusion models [15], applying pose masks to the attention module of the ViT [2]. These strategies effectively guide the network to focus on pose regions, thereby improving pose fidelity.
While precise pose alignment is essential, high-quality human image generation also demands the guidence of global visual semantics to ensure overall image quality. However, these methods [15, 1, 2] emphasize the modeling of pose details while neglecting overall image quality. To address this issue, we propose Knowledge Based Global Guidance and Dynamic pose Masking for Human Image Generation (KB-DMGen). KB-DMGen introduces a visual Knowledge Base (KB) to provide global visual semantics related to the input text during image generation, improving pose accuracy while maintaining image quality. Meanwhile, the Dynamic Masking (DM) mechanism adaptively adjusts the weights of pose-related regions, enabling the model to better balance local pose precision compared with Stable-pose [2]. By injecting them into different stages of the diffusion process, our framework enhances pose accuracy through both global and local aspects without compromising image quality. In summary, the contributions of this paper are as follows:
  • Designing a visual KB to improve pose generation accuracy while preserving image quality.
  • Designing DM to control the precise generation of poses.
  • Injecting KB and DM into different stages of the SD [12] to enhance pose accuracy via global and local control without compromising image quality.
@@
Figure 3: (a) The codebook training. $V$ is the image feature after mapping. $q(\cdot)$ denotes the nearest-neighbor mapping of $V$ in the codebook embedding space, where $y$ is the nearest neighbor index and $V_{q}$ is the image feature quantized by the codebook. (b) The process of classifier training. The text encoder is consistent with $\tau_{\theta}$ in Fig. 2. $\hat{y}$ denotes the codebook entry assignment of visual tokens predicted from text features, supervised by $y$ obtained in (a).
@@ II Related work
Pose-Guided Human Iage Generation. Previous pose-guided approaches take a source image and pose as input to generate photo-realistic human images with preserved appearance, relying on GANs/VAEs for conditional generation [7, 9, 8, 16, 17]. Recently, Bhunia et al. [18] introduce texture diffusion modules and disentangled classifier-free guidance to accurately model appearance-pose relationships and ensure input-output consistency. Zhang et al. [19] propose a novel Dual-task Pose Transformer Network (DPTN), which introduces an auxiliary task (i.e., source to source task) and exploits the dual-task correlation to promote the performance of PGPIG. Shen et al. [20] presented a Progressive Conditional Diffusion Model (PCDM), which narrows the gap between character images under target poses and source poses step by step via a three-stage procedure. The above methods [18, 19, 20] remain highly dependent on the distribution of source images, operating under the paradigm of pose transfer that necessitates original image inputs. In contrast, recent work based on pose guided T2I dispenses with the need for source images entirely: HumanSD [1] enhances the pose accuracy of HIG via heatmap-guided losses and Stable-Pose [2] employs coarse-to-fine masking for precise HIG. GRPose [15] establish a graph topological structure between the pose priors and latent representation of diffusion models to capture the intrinsic associations between different pose parts. However, these methods [1, 2, 15] prioritize pose fidelity and lack global image quality assurance, prompting us to introduce a visual knowledge base, which enhances the pose accuracy of HIG and ensures image quality via global guidance and a dynamic mask to guarantee pose precision via local control.
@@
Relation to VQ-VAE [28]. The core idea of Vector Quantised-Variational AutoEncoder (VQ-VAE)[28, 29, 30] is to learn discrete latent representations by mapping continuous features into a codebook of quantized vectors. This foundational framework has been used in various tasks. For example, PCTPose[31] uses a discrete codebook to model 2D human joint relations, while VQ-VAE-2 [32] extends VQ-VAE for large-scale image generation. These methods [32, 31] employ EMA updates to mitigate codebook collapse. However, such updates rely on additional hyperparameters (e.g., decay rates) and adapt slowly. Moreover, the pixel-wise sampling in VQVAE2 [32] is time-consuming. In addition, Zero-shot T2I [21] explores VQ-VAE-based frameworks for text-to-image generation, and Text2Human [33] uses a VQ-VAE-style codebook to model local clothing attributes in human image generation. In contrast, during the separate training of KB, our method focuses on aligning language and vision through a shared discrete token space. Specifically, we discretize image features to form a visual semantic codebook, introduce an entropy-based loss to prevent codebook collapse, and train a classifier to retrieve relevant tokens from the visual semantic codebook based on textual prompts, providing more efficient semantic guidance for image generation.
@@ III Method
This section presents the overall framework, followed by detailed codebook in the KB and classifier, concludes with the DM strategy, and finally introduces the loss of KB-DMGen.
@@ III-A Overview of KB-DMGen
Our goal is to generate high-quality human images conditioned on pose priors. To this end, we propose KB-DMGen, a Stable Diffusion based framework equipped with two adapters: a KB Adapter and a DM Adapter (Fig. 2). These modules enhance pose accuracy while preserving global image fidelity. Specifically, SD [12] provides the backbone, where an input image $x\in\mathbb{R}^{H\times W\times 3}$ is encoded into latent $z_{0}=E(x)$, perturbed into $z_{t}$ by Gaussian noise, and denoised across $T$ steps. At each step, a U-Net conditioned on text $p$, pose priors $\varphi$, KB, and DM predicts the noise to reconstruct $\hat{x}$.
@@ III-B Knowledge Base
The KB is composed of diverse codebook entries, each encoding distinct visual feature priors. The training of KB consists of two stages: first, aligning image features to the codebook vectors, and second, training a classifier to predict codebook assignments. In the final stage, the trained KB is integrated into the diffusion model to guide image generation.
@@
Stage 1: Codebook training. The goal of Codebook Training is to learn a discrete representation of image features that can be queried by text. As shown in Fig. 3a, similar to VQVAE [28], we freeze the CLIP encoder and add a projection layer to process the encoded features, facilitating stable training of the codebook. The codebook consists of trainable entries $\mathcal{Z}={(e_{1},e_{2},\dots,e_{K})}$, where $\mathcal{Z}\in\mathbb{R}^{K\times C}$, with $K$ denoting the number of codebook entries and $C$ the feature dimension.
@@
Given an image, we encode it to obtain dense image features $V=(v_{1},v_{2},\cdots,v_{N})\in\mathbb{R}^{N\times C}$, where $N=H\times W$ is the number of spatial tokens and $C$ is the feature dimension. Each token $v_{i}$, $i\in\{1,2,\cdots,N\}$, is quantized via a nearest-neighbor lookup in the codebook embedding space, as defined by the following equation:
$q(v_{i}=k|V)=\begin{cases}1&\text{if }k=\arg\min_{j}\left\|v_{i}-e_{j}\right\|_{2}\\ 0&\text{otherwise},\end{cases}$ (1)
where $j\in\{1,2,\cdots,K\}$ and $q(v_{i}=k|V)$ denotes a one-hot indicator. We use $q(v_{i})$ to represent the index to the corresponding codebook entry. We denote the set of codebook indices for image token as $y=(q(v_{1}),q(v_{2}),\cdots,q(v_{N}))$, which not only allows us to select the quantized image features $V_{q}=(e_{y_{1}},e_{y_{2}},\cdots,e_{y_{N}})\in\mathbb{R}^{N\times C}$, but also serves as the target labels for subsequent classifier training. Finally, the VQ-VAE [28] loss consists of a reconstruction term and a commitment term, formulated as:
$\mathcal{L}_{\text{VQ}}=\left\|\text{sg}[V_{q}]-V\right\|^{2}+\beta\left\|V_{q}-\text{sg}[V]\right\|^{2},$ (2)
where sg denotes stop-gradient and $\beta=0.25$.
After the basic VQ-VAE objective, we encourage balanced codebook usage. Without regularization, training often collapses to a few entries. To mitigate this, we add an entropy loss that promotes diverse token assignments. Formally, recall that the quantization process yields the index set of image tokens $y=(q(v_{1}),q(v_{2}),\cdots,q(v_{N}))$, where each $q(v_{i})$ corresponds to the selected codebook entry. To measure the usage distribution of the codebook, we define the empirical frequency of token assignments as:
$p_{k}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}[q(v_{i})=k],\quad p\in\mathbb{R}^{K},$ (3)
where $K$ denotes the number of codebook entries and $p_{k}$ is the average usage probability of the $k$-th entry. Based on this distribution, we compute the Shannon entropy:
$\mathcal{H}(p)=-\sum_{k=1}^{K}p_{k}\log(p_{k}+\epsilon),$ (4)
where $\epsilon$ is a small constant for numerical stability. The entropy is then normalized by the maximum entropy $\log K$, yielding a value in $[0,1]$:
$\mathcal{H}_{\text{norm}}(p)=\frac{\mathcal{H}(p)}{\log K}.$ (5)
Finally, the entropy regularization loss is defined as:
$\mathcal{L}_{\text{entropy}}=1-\mathcal{H}_{\text{norm}}(p),$ (6)
which penalizes skewed token usage and encourages uniform allocation across all codebook entries. The overall codebook training objective becomes:
$\mathcal{L}=\mathcal{L}_{\text{VQ}}+\mathcal{L}_{\text{entropy}}.$ (7)
@@
Stage 2: Classifier training. This stage trains a classifier to align text features with the learned discrete visual tokens (codebook entries). As shown in Fig. 3b, Given an input text, we extract a embedding feature $F$ using a freezed pretrained text encoder (e.g., CLIP). Next is the setting of the classifier. The dimension of the result after $F$ is flattened and projected is changed:
$X=\text{Linear}(\text{flatten}(F)).$ (8)
where $X\in\mathbb{R}^{N\times V}$. $N$ matches the number of image tokens, hence the same symbol and $V$ is feature dimension. subsequently, like the operation of PCTPose [31], four MLP-Mixer [34] blocks is used to process the features $X$, and output the logits of token classification:
$\hat{y}=\mathcal{M}(X),$ (9)
where $\hat{y}$ has the shape of $\mathbb{R}^{N\times K}$ and $K$ is the number of ocdebook entries. The supervision $y$ is obtained from a pretrained and frozen stage 1, which takes the input image corresponding to the current text inputs. We optimize $\hat{y}$ against $y$ using cross-entropy loss:
@@
This learning process encourages the alignment of textual features with the corresponding visual codebook entries, thus enabling text-conditioned visual token retrieval. When this stage is trained, the inference stage is shown by the yellow arrow of Fig. 3b, where $V_{q}$ is the retrieved result.
@@
KB-DMGen 53.47 72.33 1.56 10.54 2.54 32.43
Stage3: KB Embedding Diffusion Model. In this stage, the pretrained KB is integrated into a diffusion model to guide the image generation process. Fig. 2 illustrates that the text $p$ is encoded by text encoder $\tau_{\theta}$ to query the KB and generate visual semantic priors $z_{t}^{{}^{\prime\prime}}$. During stage-2 inference (Fig.3b), the feature $V_{q}$ is equivalent to $z_{t}^{{}^{\prime\prime}}$. To achieve the integration of $z_{t}^{{}^{\prime\prime}}$ and U-Net, we employ a FiLM-style [35] modulation block. Concretely, the feature $z_{t}^{{}^{\prime\prime}}$ retrieved from the codebook $\mathcal{Z}$ is passed through an Multi-Layer Perceptron (MLP) [36, 37] to generate the affine parameters $(\gamma_{cb},\beta_{cb})$. Meanwhile, the diffusion time step $t$ embedding $\tau_{t}\in\mathbb{R}^{C_{t}}$ is projected to generate $(\gamma_{t},\beta_{t})$, which serves as a global modulation to further ensure overall image quality. The final modulation parameters are obtained by combining both sources:
$\gamma=\gamma_{cb}\odot\gamma_{t},\quad\beta=\beta_{cb}\odot\beta_{t},$ (11)
where $\odot$ means element-wise multiplication. The $z_{t}^{{}^{\prime\prime\prime}}$ in Fig. 2 can be obtained and injected U-Net:
$z_{t}^{{}^{\prime\prime\prime}}=z_{t}\odot(1+\gamma)+\beta,$ (12)
where $(\gamma,\beta)=\text{MLP}(z_{t}^{{}^{\prime\prime}},\tau_{t})$ are adapted by both visual priors $z_{t}^{{}^{\prime\prime}}$ and time step $t$. Therefore, the KB-conditioned image generation $\kappa_{\theta}$ can be formulated as:
$\kappa_{\theta}(z_{t},\mathcal{Z},t)$ $=z_{t}\odot(1+\gamma)+\beta,$ (13)
where $\mathcal{Z}$ denotes the codebook set in KB.
III-C Dynamic Masking
Our method builds upon the coarse-to-fine Pose-Masked Self-Attention (PMSA) of Stable-Pose [2], which applies Gaussian-dilated pose masks to gradually refine latent representations. The difference is that we combine Dynamic Mask and PMSA (DM-PMSA).
DM-PMSA. A binary pose mask $m_{k}$ is obtained from the skeleton image, downsampled to match the latent feature $z_{t}\in\mathbb{R}^{c\times h\times w}$ (see Fig. 2), and dilated by Gaussian kernels of decreasing sizes $\{k_{1}>\cdots>k_{N}\}$—we directly adopt the optimal Gaussian kernel configuration $\{23,13\}$ with standard deviation $\sigma=3$ from [2]—to produce $m_{k_{1}},\dots,m_{k_{N}}$}. The latent feature $z_{t}$ is then processed by a sequence of $N=2$ ViT blocks, each associated with one of the Gaussian-dilated masks in a coarse-to-fine manner. Within each block, the all patch embeddings of $z_{t}$ are projected into querie $Q$, key $K$, and value $V$, and standard attention logits [38] are computed as:
$\text{dots}=\frac{QK^{\top}}{\sqrt{d}},$ (14)
where $d$ is the projected channel. Then, Attention is restricted to pose-relevant regions and dynamically modulated:
$A_{k}=\text{softmax}\left\{(1+\delta m_{k})(\text{dots}+\text{AttnMask}(m_{k})\right)\}V,$ (15)
where $\delta=\text{Sigmoid}\{\textbf{MLP}(t)\}$ is a timestep-dependent modulation factor applied only to pose regions via $\delta m_{k}$, and $\text{AttnMask}(m_{k})\in\mathbb{R}^{l\times l}$ where $l=h\times w$ is a pose-aware attention mask: entries corresponding to pose-related patches ($m_{k}=1$) are set to 0, while all other entries are assigned -$\infty$, suppressing attention outside pose regions. This coarse-to-fine progression of $N$ blocks gradually steers the latent representation to align with the target pose, while the dynamic modulation provides additional flexibility in controlling pose influence across timesteps.
We define $F_{\theta}^{\text{dyn}}$ as the DM-PMSA process, and the conditioning function $\nu_{\theta}(z_{t},\varphi,t)$ as the combination of the DM-PMSA and the pose encoder:
$\nu_{\theta}(z_{t},\varphi,t)$ $=F_{\theta}^{\text{dyn}}(z_{t},\varphi,t)+\beta_{\theta}(\varphi),$ (16)
$z_{t}^{{}^{\prime}}$ $=z_{t}+\nu_{\theta}(z_{t},\varphi,t)$ (17)
where $\varphi\in\mathbb{R}^{h\times w\times 3}$ is the input pose skeleton in Fig. 2. In this way, $\nu_{\theta}$ captures both the spatial dependencies between body parts (via $F_{\theta}^{\text{dyn}}$ ) and global pose features (via $\beta_{\theta}$), with adaptive modulation across diffusion timesteps. Following Stable-Pose, $\beta_{\theta}$ is a trainable encoder, including six convolutional layers with SiLU activation layers, downsampling the input pose image by a factor of 8. A zero-convolutional layer is added in the end.
As shown in Fig. 2, in our framework, the injection of KB and DM is not performed at the same stage. Specifically, the DM is introduced at the early stage of the first layer ($L_{inj}=\{1\}$), leveraging the rich early image features to provide fine-grained control over local pose and structural information. In contrast, the KB is injected at intermediate and later layers ($L_{inj}=\{4,7,10\}$), providing global visual knowledge guidance in the mid-to-late stages.
III-D Loss of KB-DMGen
As shown in Fig. 2, the denoising network $\epsilon_{\theta}$ adopts a UNet backbone. Let $\epsilon_{\theta}(\mathbf{z}_{t},t,p),\ t\in\{1,\cdots,T\}$, represent a T-step denoising UNet with gradients $\nabla\theta$ over a batch and input text prompt $p$. The denoising model predicts the noise as:
$\epsilon_{\text{pred}}=\epsilon_{\theta}(z_{t},t,\tau_{\theta}(p),\nu_{\theta}(z_{t},\varphi,t),\kappa_{\theta}(z_{t},\mathcal{Z},t)),$ (18)
where $\tau_{\theta}$ is the text encoder, $\nu_{\theta}(z_{t},\varphi,t)$ is DM conditioning function in Sec. III-C and $\kappa_{\theta}(z_{t},\mathcal{Z})$ is the KB-guided conditioning function in Sec. III-B.
The reconstruction error outside the pose regions is computed as:
$\mathcal{L}_{\text{um}}=\underset{\begin{subarray}{c}\mathbf{z},p,\varphi,\epsilon,\mathcal{Z}\sim\mathcal{N}(0,I),t\end{subarray}}{\mathbb{E}}\Big{[}\|(\epsilon-\epsilon_{\text{pred}})\odot(1-m_{k_{N}})\|_{2}^{2}\Big{]},$ (19)
where $m_{k_{N}}$ be the finest-level pose mask and $\epsilon$ denote the Gaussian noise added to the latent encoding $z_{t}$ at timestep $t$. This term ensures that the model maintains consistency in the background and non-pose regions. For pose-relevant areas, the error is measured by:
$\mathcal{L}_{\text{m}}=\underset{\begin{subarray}{c}\mathbf{z},p,\varphi,\epsilon,\mathcal{Z}\sim\mathcal{N}(0,I),t\end{subarray}}{\mathbb{E}}\Big{[}\|(\epsilon-\epsilon_{\text{pred}})\odot m_{k_{N}}\|_{2}^{2}\Big{]}.$ (20)
This loss explicitly enforces accurate reconstruction within the pose-constrained regions. Finally, the overall training objective is a weighted sum:
$\mathcal{L}=\mathcal{L}_{\text{um}}+\alpha\mathcal{L}_{\text{m}},$ (21)
where $\alpha=5$ is a hyperparameter emphasizing the masked (pose) regions, following the optimal design of Stable-Pose [2].
Figure 4: The visualization results of KB-DMGen in single-person scenes.
Figure 5: The visualization results of KB-DMGen in multi-person scenes.
Figure 6: Qualitative comparison. The KB (w/o $\gamma_{t}\&\beta_{t}$)+DM (w/o Sig.) combination exhibits a synergistic effect, enhancing both pose accuracy and image quality beyond individual contributions. Sig. means Sigmoid.